# It’s Morphing Time: Unleashing the Potential of Multiple LLMs via Multi-objective Optimization

### v2

@@ 5.4 Ablation Study
MOBO + Sparsity + W2S Single Model 1 + 2 + 3 58.83 72.0 52.1 57.54 53.66
@@
In this section, we conduct ablation studies on all modules of MM-MO to verify their respective effectiveness. We use the scores of the merged models obtained by the DARE-TIES without configuration search method on C-EVAL, GSM8K, and HumanEval as baselines.
1. BO (C-EVAL) & BO (GSM8K). First, we use a single-objective BO algorithm to search configurations for DARE-TIES model merging. As shown in Table 4 , BO (C-EVAL) achieves higher accuracy on C-EVAL, while BO (GSM8K) performs outstandingly on the GSM8K dataset. Additionally, HumanEval, which was not listed as an optimization target, also shows some improvement, indicating the importance and effectiveness of merging configuration searches for model merging.
@@
5. MOBO + Sparsity + W2S. On the basis of MOBO, additional sparsity metrics are introduced, and the weak-to-strong method is employed. The model merged using this method performs very well on C-EVAL, and also maintaining strong overall performance.
@@
6. MOBO + Sparsity + W2S + Fisher. This is the MM-MO method proposed in this paper, which builds on MOBO by incorporating sparsity metrics, the weak-to-strong method, and Fisher information to select candidate solutions. These three contributions ensure that the final model merging configuration found has the best overall performance. As seen in Table 4, it achieved the highest average score of 59.47, with the highest accuracy on the C-EVAL (hard), GSM8K, and HumanEval.
@@
This series of ablation studies demonstrates that the proposed improvements to the acquisition strategy (weak-to-strong method & Fisher information) and additional optimization objective (sparsity metric) all positively contribute to model merging configuration searches using MOBO. The weak-to-strong method focuses the search on promising areas within the configuration space, increasing the possibility of finding high-quality model merging configurations; incorporating Fisher information into the environment selection process can improve the accuracy of the surrogate model, increasing the chances of discovering excellent model merging configurations within a limited number of evaluations; the sparsity metric, as an additional optimization objective, helps prevent the final merged model from overfitting to the evaluated tasks, ensuring stronger overall performance and better generalization.
6 Limitation
Our experiments were limited by computational resources, restricting our method to merging models with up to 7 billion parameters. While results were promising within this range, the method’s effectiveness for larger models remains untested. Future research is needed to assess its scalability for more extensive models, which will be key to understanding the full potential of MM-MO.
Additionally, our method currently requires source models to be homologous (fine-tuned from the same pre-trained model) to ensure compatibility. This limits its use to scenarios with identical pre-trained models. Future work will explore extending our approach to merge heterogeneous models, expanding its applicability.
7 Conclusion
@@
This paper proposes a novel approach (MM-MO) for LLM merging using multi-objective black-box optimization algorithms, addressing two major challenges: the reliance on human intuition for merging configurations and the difficulty of finding optimal configurations within a limited evaluation. Our contributions include automating the model merging process, enhancing model potential by multi-objective optimization. In addition, we improve the acquisition strategy of multi-objective optimization to deal with limited evaluations, and design an effective sparsity optimization objective to enhance the model’s generalization performance across different tasks. The experimental results demonstrate that our approach achieves superior performance compared to existing methods, offering a robust, high-performing solution for integrating diverse models.
Appendix A Pseudocode of MM-MO
Output : final configuration $\boldsymbol{X}_{Final}$
1 Initialize $N$ configurations $\boldsymbol{X}_{0}$ randomly
3 for $k=0$ to $K-1$ do
8 end for
Algorithm 1 The framework of MM-MO
Input : evaluated configurations $\{\boldsymbol{X}_{E},\boldsymbol{Y}_{E}\}$, acquisition function $\boldsymbol{qEHVI}$, surrogate model $GP$
Output : next-generation configuration $\boldsymbol{X}_{New}$
1 $\boldsymbol{X}_{C}\leftarrow$ Generate candidate configurations based on $\boldsymbol{qEHVI}$ and $\{\boldsymbol{X}_{E},\boldsymbol{Y}_{E}\}$
2 $\boldsymbol{X}_{H}\leftarrow$ Select high-quality configurations based on $\{\boldsymbol{X}_{E},\boldsymbol{Y}_{E}\}$
Algorithm 2 Generate new configuration
1. Initialization Phase.
The process begins by randomly generating an initial set of $N$ configurations, denoted as $\boldsymbol{X}_{0}$, where $N$ defaults to 10.
These configurations are evaluated using the evaluation function $\boldsymbol{f}(\boldsymbol{x})$ and the sparsity metric calculating function $\boldsymbol{g}(\boldsymbol{x})$. The results are combined into an initial set of objective values, $\boldsymbol{Y}_{0}$, which includes both evaluation results and sparsity metrics.
2. Iterative Optimization Phase:
The algorithm iterates $K$ times, where $K$ defaults to 4. In each iteration, the current set of configurations $\boldsymbol{X}_{k}$ and their associated objective values $\boldsymbol{Y}_{k}$ are used to train surrogate models (Gaussian Processes, GP) for each objective, which is a commonly used surrogate model in Bayesian optimization.
Based on these surrogate models and the current evaluated configurations, a batch of new candidate configurations ${\boldsymbol{X}}_{k}^{B}$ is generated. We set the batch size to 5 in our experiment.
The newly generated configurations are added to the existing set, and their objective and sparsity metrics are evaluated. This updates both the configuration set and their objective values for the next iteration.
3. Final Selection Phase:
After completing all iterations, the configuration with the highest objective values is selected as the final configuration $\boldsymbol{X}_{Final}$, which serves as the output of the algorithm.
During the generation of new configurations, the algorithm first generates candidate configurations based on the acquisition function (e.g., qEHVI [20]) and the previously evaluated configurations. It then selects high-quality configurations based on objective values, applies a weak-to-strong adjustment method, and finally selects half of the configurations according to the Fisher information for the next generation. This approach can increase the chances of discovering excellent model merging configurations within a limited number of evaluations. For details, please refer to Section 3.3.
Appendix B Detailed Experimental Settings
B.1 Implementation Details
@@
The maximal number of generated token is set to 1024 on all datasets. Boolean parameter do_sample is set to False for greedy decoding (always choosing the most probable next token). Experiments are conducted on one NVIDIA A100 GPU.
B.2 Details of SFT and Pre-Trained Backbones of LMs
Table 5 shows the versions and correspondences with pre-trained backbones of SFT LMs.
Table 5: Versions and correspondences with pre-trained backbones of SFT LMs.
Parameter Scale SFT LMs URL Pre-Trained Backbones
B.3 Descriptions of Datasets
In this section, we describe the datasets used in our study, each serving a specific purpose in evaluating various model abilities.
B.4 Descriptions of Model Merging Methods
We experiment with five model merging methods:
Appendix C Additional Experiments
C.1 Comparison on C-EVAL Dataset
Figure 3: Performance of Single Model and Merged Model (MM-MO) on C-EVAL.
@@ C.1 Comparison on C-EVAL Dataset
Figure 3 visualizes the accuracy of single models and the MM-MO merged model in answering four different categories of C-EVAL questions, the average C-EVAL accuracy, and the accuracy on the hard C-EVAL questions. As shown in the figure, MM-MO (red) covers the largest area and demonstrates superior performance. Additionally, in all four categories of C-EVAL tasks—STEM, Social Sciences, Humanities, and Others—the MM-MO merged model achieves higher accuracy than any single best-performing model, which intuitively demonstrates that using MM-MO for model merging can combine the strengths of various models and unlock their potential.
C.2 Case Study: Mathematical Reasoning
Table 6: Case Study on GSM8K.
@@ C.2 Case Study: Mathematical Reasoning
Table 6 provides a comparative analysis of responses to a mathematical reasoning problem by three models. By employing MM-MO to search for the optimal configuration for model merging, we enhanced our model’s performance in mathematical reasoning capabilities.
@@
Specifically, in this problem, only our merged model gives the correct answer. Our model (merged by MM-MO method) understood the question well and followed the step by step instructions to give the final correct answer. However, the model merged by TIES method, while also following the step by step instructions, does not properly understand the problem and give a mathematical reasoning process that meets our requirements. Finally, the model merged by DARE-TIES method fails entirely, providing irrelevant and repetitive responses, despite DARE-TIES utilizing a dropout-like method to avoid parameter conflicts during model merging, which generally results in better performance compared to TIES, it can suffer from significant performance drops on specific problems if the model merging configuration is not correctly set. This issue is evident in the provided case study, where DARE-TIES fails to deliver a coherent response. This highlights the importance of correctly configuring model merges to retain crucial parameters, ensuring consistent performance across diverse problem sets. The MM-MO model’s success demonstrates the effectiveness of our model merging strategy in overcoming these challenges and providing accurate solutions.
C.3 Parameter Sensitivity Experiment
Table 7: Performance of MM-MO with different number of function evaluations.
Single Model 1 Qwen1.5-7B-Chat 55.34 68.7 51.1 54.59 46.95
Single Model 2 Liberated-Qwen1.5-7B 55.47 69.7 50.1 53.30 48.78
Single Model 3 firefly-qwen1.5-en-7b 51.01 70.0 50.7 49.81 33.54
MM-MO / FE = 20 Single Model 1 + 2 + 3 54.62 70.7 51.5 53.60 42.68
MM-MO / FE = 40 Single Model 1 + 2 + 3 59.27 72.0 52.8 57.62 54.66
MM-MO / FE = 30 (Default) Single Model 1 + 2 + 3 59.47 71.9 52.7 57.77 55.49
In this section, we compare the performance of MM-MO with different number of function evaluations to validate the sensitivity of our proposed method to this parameter. Table 7 illustrates the impact of varying number of function evaluations on the performance of MM-MO.
C.4 Time Overhead of MM-MO
Table 8: Time overhead of MM-MO for different model size.
Merging Method Model Size Time Overhead (min)
MM-MO 1.5B 259.21
MM-MO 7B 526.52
Table 8 lists the time overhead of MM-MO for different model size, 1.5B and 7B. Compared to fine-tuning on large datasets to improve the overall performance of the model, using MM-MO for model merging with existing models does not require additional training data. It only needs to be evaluated on the existing language model evaluation datasets to quickly obtain a merged model with great overall performance in a relatively short period of time.
C.5 The Final Configuration Searched using MM-MO
Figure 4: Final Configuration Searched using MM-MO. (1.5B & 7B)
Figure 4 illustrates the final parameter configuration of model merging (1.5B & 7B) searched using MM-MO. Density is the fraction of weights in differences from the base model to retain and weight represents the relative weighting of a given model. From the figure we can know the contribution of different single models to the final model merging.
Appendix D Licenses
The licenses for the codes we used in this work are shown in Table 9.
@@ Appendix D Licenses
Table 9: List of licenses for the codes suite we used in this work.

### v3

@@ VI-G Ablation Study
MOBO + FI Single Model 1 + 2 + 3 58.55 71.8 52.6 56.97 52.83
@@
As shown in Table X, we conduct ablation studies on all modules of MM-MO to verify their respective effectiveness. The evaluation focuses on the scores of the merged models across three benchmarks: C-EVAL, GSM8K, and HumanEval, which respectively assess the models’ overall comprehensive ability, mathematical reasoning skills, and coding capabilities. Additionally, C-EVAL (hard) is a subset of particularly challenging subjects within C-EVAL, designed to test advanced reasoning abilities and problem-solving skills.
1. BO (C-EVAL) & BO (GSM8K). First, we use a single-objective BO algorithm to search configurations for DARE-TIES model merging. As shown in Table X , BO (C-EVAL) achieves higher accuracy on C-EVAL, while BO (GSM8K) performs outstandingly on the GSM8K dataset. Additionally, HumanEval, which was not listed as an optimization target, also shows some improvement, indicating the importance and effectiveness of merging configuration searches for model merging.
@@
5. MOBO + FI. The acquisition strategy in MOBO is improved by introducing the environment selection based on Fisher information (FI), which can enhances the accuracy of the surrogate model more rapidly, and increases the possibility of discovering superior model merging configurations. Experimental results demonstrate that the MOBO + FI method achieves better model merging configurations compared to MOBO alone, given the same number of evaluations.
@@
6. MOBO + Sparsity + W2S + FI. This is the MM-MO method proposed in this paper, which builds on MOBO by incorporating sparsity metrics, the weak-to-strong method, and Fisher information to select candidate solutions. These three contributions ensure that the final model merging configuration found has the best overall performance. As seen in Table X, it achieved the highest average score of 59.47, with the highest accuracy on the C-EVAL, GSM8K, and HumanEval.
@@
This series of ablation studies demonstrates that the proposed improvements to the acquisition strategy (weak-to-strong method & Fisher information) and additional optimization objective (sparsity metric) all positively contribute to merging configuration searches using MOBO. The weak-to-strong method focuses the search on promising areas within the configuration space, increasing the possibility of finding high-quality model merging configurations; incorporating Fisher information into the environment selection process can improve the accuracy of the surrogate model, increasing the chances of discovering excellent model merging configurations within limited number of evaluations; the sparsity metric, as an additional optimization objective, helps prevent the final merged model from overfitting to the evaluated tasks, ensuring stronger overall performance and better generalization.
VI-H The Final Configuration Searched via MM-MO
Figure 6: Final Configuration Searched via MM-MO. (Backbone: Qwen1.5-7B & Qwen2-1.5B)
Fig. 6 illustrates the final parameter configuration of model merging (1.5B & 7B) searched using MM-MO. Density is the fraction of weights in differences from the source model to retain and weight represents the relative weighting of a given model. From the figure we can know the contribution of different single models to the final merged model.
Taking the 7B models as an example, the higher Density and Weight values of Qwen1.5-7B-Chat and Liberated-Qwen1.5-7B suggest that these two models play a critical role in shaping the merged model. In contrast, the relatively lower values of firefly-qwen1.5-en-7b indicate that its influence is minimal. This result aligns with the scores of these three models presented in Table III, showing the efficacy and substantial promise of MM-MO for model merging.
VII Conclusion and Future Work
@@
In this paper, we introduced a novel framework that formalizes model merging configuration search as a multi-objective optimization problem. Building on this framework, we proposed a novel approach MM-MO for LLM merging using multi-objective black-box optimization algorithms, addressing two major challenges: the reliance on human intuition for merging configurations and the difficulty of finding optimal configurations within limited evaluations. Our contributions include automating the model merging process, enhancing model potential by multi-objective optimization. In addition, we improve the acquisition strategy of multi-objective optimization to deal with limited evaluations, and design an effective sparsity optimization objective to enhance the model’s generalization performance across different tasks. The experimental results demonstrate that our approach achieves superior performance compared to existing methods, offering an automatic and efficient solution for integrating diverse models.
Our method currently requires source models to be homologous (fine-tuned from the same pre-trained model) to ensure compatibility, which is also the mainstream approach in current model merging studies [27, 17]. Future efforts will made to extend our approach to merge heterogeneous models, thereby expanding its applicability. Additionally, our experiments were limited by computational resources, restricting us to merging models up to 13 billion parameters. While results were promising within this range, the method’s effectiveness for larger models remains untested. Future research is needed to assess its scalability for more extensive models, which will be key to understanding the full potential of MM-MO.
Appendix A Detailed Experimental Settings
A-A Implementation Details
@@
The maximal number of generated token is set to 1024 on all datasets. Boolean parameter ’do_sample’ is set to ’False’ for greedy decoding (always choosing the most probable next token). Experiments are conducted on one NVIDIA A100 GPU.
Appendix B Additional Experiments
B-A Comparison on C-EVAL Dataset
Figure 7: Performance of single and merged LLMs on C-EVAL.
@@ B-A Comparison on C-EVAL Dataset
Figure 7 visualizes the accuracy of single models and the MM-MO merged model in answering four different categories of C-EVAL [31] questions, the average C-EVAL accuracy, and the accuracy on the hard C-EVAL questions. As shown in the figure, MM-MO (red) covers the largest area and demonstrates superior performance. Additionally, in all four categories of C-EVAL tasks—STEM, Social Sciences, Humanities, and Others—the MM-MO merged model achieves higher accuracy than any single best-performing model, which intuitively demonstrates that using MM-MO for model merging can combine the strengths of various models and unlock their potential.
B-B Case Study: Mathematical Reasoning
TABLE XI: Case Study on GSM8K.
@@ B-B Case Study: Mathematical Reasoning
Table XI provides a comparative analysis of responses to a mathematical reasoning problem by three models. By employing MM-MO to search for the optimal configuration for model merging, we enhanced our model’s performance in mathematical reasoning capabilities.
@@
Specifically, in this problem, only our merged model gives the correct answer. Our model (merged by MM-MO method) understood the question well and followed the step by step instructions to give the final correct answer. However, the model merged by TIES method [27], while also following the step by step instructions, does not properly understand the problem and give a mathematical reasoning process that meets our requirements. Finally, the model merged by DARE-TIES method [17] fails entirely, providing irrelevant and repetitive responses, despite DARE-TIES utilizing a dropout-like method to avoid parameter conflicts during model merging, which generally results in better performance compared to TIES, it can suffer from significant performance drops on specific problems if the model merging configuration is not correctly set. This issue is evident in the provided case study, where DARE-TIES fails to deliver a coherent response. This highlights the importance of correctly configuring model merges to retain crucial parameters, ensuring consistent performance across diverse problem sets. The MM-MO model’s success demonstrates the effectiveness of our model merging strategy in overcoming these challenges and providing accurate solutions.
B-C Case Study: Coding
TABLE XII: Case Study on MBPP.
Problem
Instruction You are an expert Python programmer, and here is your task: Write a python function to find the cube sum of first n natural numbers. Your code should pass these tests: assert sum_Of_Series(5) == 225 assert sum_Of_Series(2) == 9 assert sum_Of_Series(3) == 36
Ground Truth Answer 1 def sum_Of_Series(n): # Initialize sum cube_sum = 0 # Iterate through numbers from 1 to n for i in range(1, n + 1): cube_sum += i ** 3 return cube_sum
Ground Truth Answer 2 def sum_Of_Series(n): # Formula for the sum of cubes of first n natural numbers return (n * (n + 1) // 2) ** 2
Answers
MM-MO (Ours) [BEGIN] ’def sum_Of_Series(n): sum = 0 for i in range(1,n+1): sum += i**3 return sum’ [DONE]
TIES [BEGIN] ’def sum_Of_Series(n): return n*(n.n*(n.n+1))//2’ [DONE]
DARE-TIES [BEGIN] ’def sum_Of_Series(n): return n*(n*(n+1)*(n+2))//6’ [DONE]
Table XII illustrates a case study of the merged models derived from different methods in addressing a code-related problem. In this case, the task was to generate a Python function to compute the cube sum of the first n natural numbers. Although this is a relatively simple problem, only the merged model obtained through the MM-MO method provided a completely correct answer, accurately generating the desired function using an iterative approach.
In contrast, the merged models obtained via the TIE and DARE-TIES methods adopted a formula-based approach to implement the function. However, the TIE-based model introduced an invalid syntax, specifically n.n, likely due to parameter conflicts during the model merging process. Meanwhile, the DARE-TIES-based model employed an incorrect cube sum formula.
This case clearly demonstrates the superior stability and performance of the merged model obtained through the MM-MO configuration search.
B-D MM-MO with More Source Models
TABLE XIII: Performance comparison of different merging methods and single models.
Merging Method Models C-EVAL MMLU GSM8K Human Eval MBPP
Single Model 1 Qwen1.5-7B-Chat 68.7 60.06 54.59 46.95 34.20
Single Model 2 Liberated-Qwen1.5-7B 69.7 58.84 53.30 48.78 38.80
Single Model 3 firefly-qwen1.5-en-7b 70.0 51.66 49.81 33.54 28.40
Single Model 4 TechGPT-2.0 69.4 60.13 57.85 47.56 39.80
DARE + TIES Single Model 1 + 2 + 3 + 4 68.1 59.70 62.02 47.56 35.80
DARE + TIES w/ EMM Single Model 1 + 2 + 3 + 4 67.8 60.12 63.00 38.41 27.40
DARE + TIES w/ MM-MO (Ours) Single Model 1 + 2 + 3 + 4 69.5 60.53 66.03 49.39 39.80
In this section, we increased the number of models used for merging to four, aiming to evaluate the performance of MM-MO when merging a larger number of models. Specifically, in addition to the three source models outlined in the experimental setup of main paper, Qwen1.5-7B-Chat [3], Liberated-Qwen1.5-7B [39], and firefly-qwen1.5-en-7b [40], we incorporated TechGPT-2.0 [wang2024techgpt], which was trained on a large-scale Chinese academic corpus.
As shown in Table 3, we compared the performance of the source models and the merged model across five different tasks. Following Qwen [3], the test questions span four domains: Chinese, English, Mathematics, and Coding. The experimental results demonstrate that MM-MO maintains excellent performance even as the number of source models increases. Using our proposed MM-MO method, we were able to search for a better model merging configuration, enabling the merged model to outperform the baselines across all tasks.
Furthermore, it is noteworthy that the merged model obtained through MM-MO surpassed the performance of even the best individual source model across all five tasks. This indicates that the MM-MO method effectively integrates the strengths of different source models, resulting in a unified, high-performing model.
B-E MM-MO with Validation Set in Different Languages
TABLE XIV: Performance comparison of MM-MO with Validation Set in Different Languages. (Backbone: Qwen1.5-7B)
Merging Method C-EVAL MMLU GSM8K Human Eval MBPP
DARE + TIES 69.5 60.06 55.72 49.39 37.80
DARE + TIES w/ MM-MO (MMLU) 69.6 60.48 58.38 51.83 39.00
DARE + TIES w/ MM-MO (C-EVAL) 71.9 60.81 57.77 55.49 42.20
In MM-MO, we use three objectives for optimization: the model’s score on a comprehensive evaluation set, its score on a mathematical reasoning set, and our proposed sparsity metric. The GSM8K [32] dataset is employed as the validation set to evaluate the model’s mathematical reasoning capabilities. For assessing the model’s general abilities on a comprehensive evaluation set, we utilize two different benchmarks: C-EVAL [31] and MMLU [47]. Specifically, for Qwen-series models, given their strong proficiency in Chinese, we adopt C-EVAL as the comprehensive evaluation set. In contrast, for Llama-based models, MMLU is used in the experiments.
In this section, we compare the performance differences caused by using comprehensive validation sets in different languages when merging models via MM-MO method. The backbone of merged model is Qwen1.5-7B. As shown in Table 4, although Qwen is more proficient in answering Chinese questions, using MMLU as the optimization target during the configuration search still achieves favorable results. The merged model obtained through MM-MO (MMLU) outperforms our baseline, DARE + TIES, across five different tasks, and even achieves superior performance on the GSM8K task.
Moreover, when using MM-MO (C-EVAL)—our default setting—the merged model achieves the best results in most tasks. These experimental findings suggest that MM-MO demonstrates robust performance regardless of whether C-EVAL or MMLU is used as the comprehensive validation set, further highlighting the flexibility and effectiveness of the MM-MO method in achieving high-quality model merging outcomes.
Appendix C Licenses
The licenses for the codes we used in this work are shown in Table XV.
@@ Appendix C Licenses
TABLE XV: List of licenses for the codes suite we used in this work.